# Understanding and Mitigating the Impacts of Differentially Private Census Data on State Level Redistricting

### v1

@@ 3.2 Ensemble Analysis
Markov Chain Monte Carlo (MCMC) methods enable sampling ensembles of thousands or millions of redistricting plans from predefined distributions for analysis [17, 18, 19]. Our paper makes use of the $\mathop{\sf ReCom}$ algorithm, a merge-split MCMC algorithm which has been used by expert witnesses in gerrymandering cases to generate ensembles [20, Harper v. Hall 2022, Allen v. Milligan 2023]. These techniques are flexible, allowing for many constraints to be taken into account in the ensemble generation process. For example, [21] illustrate how ensembles can be used in the enforcement of the Voting Rights Act.
We use the implementation of the $\mathop{\sf ReCom}$ algorithm in the GerryChain software package [22]. In Section 6 we modify $\mathop{\sf ReCom}$ to select plans with more majority-minority districts (MMDs) using an optimization technique called short bursts, which has been shown to outperform other techniques including biased random walks and simulated annealing [23]. We split our short burst chains into a series of 10 sub-chains, each starting from the same initial partition. This was suggested to us by one of the authors of [20], and we found that it helped in discovering plans with more majority-minority districts.
Unless otherwise noted, we require population deviation at most 5% from the ideal, use block groups as the smallest unit of geography, and subsample our ensemble from a run of 1,000,000 steps. Using the Gelman-Rubin split-$\hat{R}$ diagnostic [24], we find that estimates of the metrics we study in Section 5 show clear signs of convergence in the large majority of the geographies we study ($\hat{R}\leq 1.01$, $\mathsf{ESS}\geq 400$). We see clear signs of convergence in a smaller majority ensembles in Section 6. In that context, we view the convergence tests as less important: we are using MCMC sampling more as a way to optimize quantity of interest (the number of MMDs) rather than to estimate one (the MMD discrepancy), in line with the Gingles 1 test. See Table 3 in the Supporting Information for more information on chain convergence tests.
We generate ensembles for each of 93 state legislative geographies. This includes the upper and lower geographies for 45 bicameral states and one geography for Nebraska’s unicameral legislature. It excludes 4 states where our MCMC chains do not find any plans that are population balanced to 5% in DEMO: Hawaii, New Hampshire, North Dakota, and Vermont. This limitation is due to our use of Census block groups rather than Census blocks as the unit of geography in our chains. Since the state legislative districts in these states are relatively small in population, it is computationally inefficient to find valid districts composed of block groups using MCMC methods.
4 Methods in brief and prior work
@@
This paper employs ensemble analyses: we algorithmically sample large collections of possible districting plans and observe how statistics of our ensembles differ between our two data sources (see Section 3 for details). In the last five years, ensemble methods for studying gerrymandering and redistricting have made their way from academia [25, 26, 27, 28, 29] to the Supreme Court (Gill v. Whitford (2018), Rucho v. Common Cause (2019), Harper v. Moore (2022), Allen v. Milligan (2023)). [^4] Ensemble-based evidence has been more persuasive at trial courts than at the Supreme Court. It was received positively by the dissent in Rucho v. Common Cause, but less so in Allen v. Milligan: “[C]ourts should exercise caution before treating results produced by algorithms as all but dispositive of a §2 claim.”
To study the effect of the DAS, one would ideally compare the data with and without disclosure avoidance, namely the 2020 CEF and 2020 Redistricting Data. However, the CEF is confidential to those outside of the Census Bureau. We instead use two Census datasets which we call SWAP and DEMO as stand-ins for the 2020 CEF and Redistricting Data, respectively (see Section 3). The Census Bureau created DEMO created to help stakeholders study the impacts of the 2020 DAS by comparing it to SWAP. Both SWAP and DEMO are derived from the 2010 CEF, using the 2010 DAS and 2020 DAS, respectively. See Section 7.1 for further discussion on the limitations of these datasets.
At a very high level, our approach is simple. First, we generate an ensemble of 100,000 plans, say. These plans are drawn to satisfy some constraint—or to maximize some objective—on the DEMO dataset (standing in for 2020 Redistricting Data). Second, across plans in the ensemble, we see how often the constraint is violated—or how the value of the objective compares—on the SWAP dataset (standing in for 2020 CEF).
We adopt our basic setup from Kenny et al. [6], who also study the effects of the 2020 DAS on redistricting through an ensemble-based comparison of DEMO and SWAP. They show that a significant fraction of redistricting plans satisfying a given maximum population deviation $\tau$ on DEMO exceed the $\tau$ limit on SWAP. For example, for $\tau=1\%$, about $75\%$ of plans exceed the deviation limit on SWAP; for $\tau=5\%$, about 5% of plans do [6, Figure S4.4]. Kenny et al. [6] conclude from this that “the added noise makes it impossible to follow the principle of One Person, One Vote, as it is currently interpreted by courts and policy-makers,” pointing to state legislative districts as particularly challenging in this regard.
We disagree. In Section 5, we propose a simple remedy, while also greatly extending the above analysis—using 93 state legislative geographies (instead of 1), and ensembles of 100,000 plans (instead of 5,000). Kenny et al. [6] also briefly consider how the DAS affects the apparent number of majority-minority districts that can be drawn—the subject of Section 6 of this paper. Contrary to our results, they observe fewer majority-minority districts in DEMO versus SWAP. We suspect this is because their analysis is based on inferred race—using a technique called BISG on records of registered voters—rather than the race data in the redistricting files themselves. [^5] Other results from the same case study suffice for the effect they observe. Comparing DEMO and SWAP, the BISG-inferred “proportions of Black and Hispanic [registered voters] are much smaller, especially in the blocks where they form a majority group.” We more closely follow the Gingles 1 test by using the race data in the Redistricting Data. See Section 3 for more details on ensemble generation.
@@ 4.1 Notation
Let $\mathsf{data}$ denote a reference dataset, either DEMO or SWAP. A state-legislative redistricting plan $\mathcal{P}$ is a partition of a state into $k$ districts: $D_{1}$, …, $D_{k}$. Each district $D_{i}$ is a contiguous collection of census blocks. We denote a district’s population in $\mathsf{data}$ as $\mathsf{pop}_{\mathsf{data}}(D)$, and it’s voting-age population $\mathsf{VAP}_{\mathsf{data}}(D)$.
@@ 5 Balancing district populations with noise
A significant fraction of redistricting plans satisfying a given population deviation limit on DEMO exceeds that limit on SWAP, as first shown by Kenny et al. [6]. They conclude that as a result it is “impossible to follow the principle of One Person, One Vote,” pointing to state legislative districts as particularly challenging in this regard.
This section challenges that conclusion. We propose a straightforward method for ensuring population balance under the unseen dataset, specifically focusing on state legislative redistricting. Our method—using a small tightening of the acceptable population deviation threshold—is very effective, and does not greatly increase the difficulty of finding valid plans nor require the adoption of new redistricting techniques.
As noted in Section 2.1, the lack of a de minimis population deviation threshold in Congressional redistricting belies the uncertainty inherent in census tabulations; we can only speculate on whether the Court would deem privacy-induced uncertainty exceptional in that regard. Instead, we focus on state legislative districts, where tolerances are looser but precise thresholds are still observed. See Section 7 for a discussion of further limitations, especially the impact of using block groups as our geounit for redistricting.
@@ 5.1 Meeting population deviation limits using offsets
In this section, we are concerned with data user’s ability to use DEMO to generate redistricting plans satisfying a maximum population deviation of $\tau$ under SWAP (equivalently, under the CEF). We propose a simple approach: draw the plan to satisfy a slightly tighter limit $\tau-\Delta$, for $0\leq\Delta\leq\tau$. We call $\Delta$ the offset. To evaluate this approach, we measure the discrepancy at $\tau$ with offset $\Delta$: the fraction of plans that exceed $\tau$ deviation under SWAP among an ensemble of plans with deviation at most $\tau-\Delta$ under DEMO. When $\Delta=0$, we call this the discrepancy at $\tau$ or the no-offset discrepancy rate. Finally, we also consider what we call the critical offset: the smallest offset with discrepancy less than 2% (see Appendix A for more information on the computation of the critical offset) .
5.1.1 A case study: the LA state senate
@@
Figure 1: Discrepancy with offsets in the Louisiana state senate. We plot the fraction of plans exceeding intended population tolerance limit $\tau$, with various offsets $\Delta$ for the Louisiana state senate (i.e., discrepancy at $\tau$ with offset $\Delta$). Dots (solid lines) are computed using the DEMO and SWAP datasets with ensembles of 100,000 plans for each $\tau$ and $\Delta$ (see Sec. 5.1). Squares (dashed lines) are computed from 100,000 samples from the statistical model of district populations and disclosure avoidance noise described in Sec. 5.2.
@@
We begin with a case study of the Louisiana state senate. The LA senate was used by Kenny et al. [6, Figure S4.4] to demonstrate the sort of population discrepancies we study. They measured the no-offset discrepancy for each of six population deviation thresholds $\tau$ between 0.1% and 30%. We replicate the analysis, summarized by the solid blue line in Figure 1. The overall conclusion remains the same: many plans drawn to a satisfy $\tau$ population deviation using DEMO exceed that threshold on SWAP. The exact numbers differ, perhaps because we use block groups as the building block (not precincts), a different MCMC sampler implementation, and larger ensembles.
@@
To test the effectiveness of offsets as a mitigation, we measure the discrepancy at $\tau$ with twenty values of the offset parameter $\Delta$ between 0% to 1%, for the same six values of $\tau$ as before. The solid lines in Figure 1 illustrate the results for $\Delta=0.05\%$, 0.1%, and 0.2%. Using $\Delta=0.1\%$ reduces the discrepancy at $\tau=5\%$ from 18.6% to 1.2%. Using $\Delta=0.2\%$ reduces the discrepancy at $\tau=1\%$ from 59.1% to 0.1%. Notably, the discrepancy at $\tau=5\%$ with offset $\Delta=0.2\%$ was 0. That is, of 100,000 plans generated using a 4.8% tolerance on DEMO, all satisfied a 5% tolerance on SWAP.

### v2

@@ 3.2 Ensemble Analysis
In the past decade, algorithmic methods for studying gerrymandering and redistricting have made their way from academia [18, 11, 9, 16, 20] to the Supreme Court (Gill v. Whitford (2018), Rucho v. Common Cause (2019), Harper v. Moore (2022), Allen v. Milligan (2023)). [^4] Ensemble-based evidence has been more persuasive at trial courts than at the Supreme Court. It was received positively by the dissent in Rucho v. Common Cause, but less so in Allen v. Milligan: “[C]ourts should exercise caution before treating results produced by algorithms as all but dispositive of a §2 claim.”
Markov Chain Monte Carlo (MCMC) methods enable sampling ensembles of thousands or millions of redistricting plans from predefined distributions for analysis [14, 3, 19]. Our paper makes use of the $\mathsf{ReCom}$ algorithm, a merge-split MCMC algorithm which has been used by expert witnesses in gerrymandering cases to generate ensembles [13, Harper v. Hall 2022, Allen v. Milligan 2023]. We use the implementation from the GerryChain software package [36] (with some modifications, as described in Section 6).
We use voting precincts as the smallest unit of geography for drawing districts. [^5] We use the term precinct instead of the more correct “voting district” or VTD to avoid confusion with the much larger state and Congressional districts that are the main subject of this paper [39]. Keeping precincts intact is often considered desirable in practice by state legislatures and previous academic work on redistricting ensembles has used precincts to compose districts [24]. For a few states, precinct data was incompatible with our experimental setup (either due to especially large precincts or incomplete geographical data). In those cases, we used block groups as the building block of districts in our analyses. See Section 7 for a discussion of how these choices may have affected our results.
Details for population balance (Section 5)
Except where otherwise noted, each ensemble contains 100,000 plans generated by running $\mathsf{ReCom}$ for 1,000,000 steps and selecting every tenth plan. We impose various population deviation limits on the MCMC sampler, as described in Section 5.
We generate ensembles for each of 93 state legislative geographies. These include the upper and lower geographies for 45 bicameral states and one geography for Nebraska’s unicameral legislature. It excludes lower houses in 4 states (Hawaii, New Hampshire, North Dakota, and Vermont) and upper houses in 2 states (Hawaii and North Dakota) where our MCMC chains do not find any plans that are population balanced to within 5% in DEMO. This limitation is due to our use of voting precincts rather than Census blocks as the unit of geography in our chains. Since the state legislative districts in these states are relatively small in population, it is computationally inefficient to find valid districts composed of precincts using MCMC methods.
Details for majority-minority districts (Section 6)
In Section 6, we generate plans using both $\mathsf{ReCom}$ and a modified version of $\mathsf{ReCom}$ optimized to select plans with more majority-minority districts (MMDs). The latter uses a technique called short bursts, which has been shown to outperform other techniques including biased random walks and simulated annealing [8]. Details are deferred to Section 6.1.
Ensembles using unmodified $\mathsf{ReCom}$ contain 100,000 plans generated as above. Ensembles using short-bursts optimization contain 1,000 plans each (except for the Georgia state house ensemble which contains 5,000 plans).
Of the 93 state legislative geographies described above, our ensembles found at least one majority-Black district in 52 of them (27 lower house, 25 upper house). We restrict our analysis of MMD redistricting to these 52 geographies. For each geography, we impose a population deviation limit of 5% minus the critical offset found in Section 5.
Convergence
Table 3 in the Appendix summarizes results for Markov chain convergence tests for our main ensemble analyses. Using the Gelman-Rubin split-$\hat{R}$ diagnostic [22], we find that estimates of the metrics we study in Section 5 show clear signs of convergence in the large majority of the geographies we study ($\hat{R}\leq 1.01$, $\mathsf{ESS}\geq 400$). Compared to the analyses in Section 5, fewer ensembles in Section 6 exhibit clear signs of convergence. In that context, we view the convergence tests as less important: we are using MCMC sampling more as a way to optimize quantity of interest (the number of MMDs) rather than to estimate one (the MMD discrepancy), in line with the Gingles 1 test.
@@ 3.3 Notation
Let $\mathsf{data}$ denote a reference dataset, either DEMO or SWAP. A state legislative redistricting plan $\mathcal{P}$ is a partition of a state into $k$ districts: $D_{1}$, …, $D_{k}$. Each district $D_{i}$ is a contiguous collection of census blocks. [^6] In our main experiments, districts are constructed from larger units called precincts (each precinct aggregates many census blocks); we use “blocks” here only to align with the Census geographic hierarchy. See Section 7 for discussion of how using precincts (rather than blocks) may affect the observed impact of DAS noise. We denote a district’s population in $\mathsf{data}$ as $\mathsf{pop}_{\mathsf{data}}(D)$, and its voting-age population $\mathsf{VAP}_{\mathsf{data}}(D)$.
@@
4 Related Work
Two prior studies have used ensemble methods to study the effects of the 2020 DAS on redistricting, specifically on population balance and the analysis of racial gerrymandering [24, 12]. We discuss them below. Other research on the 2020 DAS include studies of the impacts on different policy issues, including minority representation [10], misallocation of federal funds [38], and public health monitoring [29], for example.
The Census Bureau has also published its own analyses of the 2020 DAS. The most relevant examines the impact of the 2020 DAS on the reliability and consistency of measurements of demographic characteristics in various geographies [48]. Among other findings, it shows that the 2020 DAS changes the apparent fraction of the population belonging to the largest demographic group by at most 5% in 95% of block groups with total population 450 to 499. Such reliability guarantees are important for Voting Rights Act enforcement, but don’t do away with issues presented by sharp thresholds, as is the focus of our work.
4.1 Kenny et al. [24]
Kenny et al. [24] study the effects of the 2020 DAS on redistricting through a variety of case studies using ensemble-based comparisons of DEMO and SWAP. [^8] Only a non-final version of DEMO (called “DAS-12.2 data” in [24]) was available at the time of their initial experiments. Results using the more recent DEMO data (“DAS-19.61 data”) are in [24, Supplementary Materials]. We adopt their basic setup: generate many maps using the DEMO data, and measure their characteristics on the SWAP data.
Population balance
Kenny et al. study of the impact on population balance for state legislative districts, with the Louisiana state senate as a case study [24]. They generate ensembles of 5,000 plans using a merge-split MCMC algorithm (the same algorithm we use, but a different implementation [27]). They find that a significant fraction of redistricting plans satisfying a given maximum population deviation $\tau$ on DEMO exceed the $\tau$ limit on SWAP [24, Figure S4.4]. For example, for $\tau=1\%$, about $75\%$ of plans exceed the deviation limit on SWAP; for $\tau=5\%$, about 5% of plans do. We include these results in Figure 1.
In Section 5, we greatly extend this analysis. We extend it in scale: using 93 state legislative geographies (instead of 1) and ensembles of 100,000 plans (instead of 5,000). We also extend it in scope: introducing offsets to mitgate discrepancies between DEMO and SWAP and measuring their effect and also proposing a model to explain the empirical observations.
Majority-minority districts
Kenny et al. [24] consider how the DAS affects the apparent number of majority-minority districts that can be drawn [24, Table S4.2], through a case study reexamining NAACP of Spring Valley v. East Ramapo Central School District (2020). They sample 10,000 plans (5% max deviation) for the school district using DEMO, and count the number of MMDs in each plan based on DEMO and SWAP. The find fewer majority-minority districts in DEMO versus SWAP.
We find the opposite, across many state legislative geographies (Section 6). As we explain next, we believe the disagreement stems from differences in how our two works (i) define and count MMDs, and (ii) generate ensembles of redistricting plans.
We define an MMD as a district having more than 50% of voting age population in the minority group. This is the measure used in Supreme Court cases including Allen v. Milligan (2023), Bartlett v. Strickland (2009), and others. Kenny et al. instead use registered voters as was done in the Ramapo case. Race data for registered voters in Ramapo are not available. To get around this, Kenny et al. count MMDs using inferred race derived from the surnames and addresses of registered voters using a technique called Bayesian Improved Surname Geocoding (BISG). BISG predicts the race or ethnicity of the registered voters in a district based on their surname and the demographic makeup of the district as measured using census data (i.e., DEMO or SWAP). Because of this, the effect of the 2020 DAS on counting MMDs as measured in [24] is mediated through its effect on BISG. The latter is a significant confounder: comparing DEMO and SWAP, the inferred “proportions of Black and Hispanic [registered voters] are much smaller, especially in the blocks where they form a majority group” [24].
The second difference is our use of an optimization technique designed to sample plans with many majority-Black MMDs. Compared to the race-blind sampler used in [24] and in Section 5, our MMD-optimizing sampler appears to exacerbate the discrepancies we observe.
4.2 Cohen et al. [12]
Cohen et al. [12] also use ensemble methods to study the effects of the 2020 DAS on redistricting. (But as we elaborate below, the methods used in that paper make it less directly relevant to the present work than it may seem.) Like [24], Cohen et al. offer a detailed study of a small handful of case studies. Our study of offsets in Section 5 was in part inspired by the proposal in [12] to tweak a standard method for measuring racially-polarized voting (the 2nd and 3rd Gingles preconditions).
While it goes a long way to understanding the scale and characteristics of the noise introduced by disclosure avoidance, the prior work does not study the interaction between the noise from disclosure avoidance and the sharp thresholds. Perhaps this is one reason that [12] and [24] arrive at opposite conclusions about whether the official Redistricting Data should be treated as ground truth for legal tests.
Finally, Cohen et al. overlook the possibility that redistricting with the perturbed data might induce additional effects. Whereas our MMD analysis creates plans using the perturbed data, the closest case study in Cohen et al. [12] only examines the demographic makeup of four existing districts in Navajo County, Arizona.
Methodology
Cohen et al. do not use the Census demonstration data whatsoever—unlike the majority of independent analyses of the 2020 DAS’s effects, including [24] and ours. Instead, they analyze samples of noised data generated using an early version of the 2020 DAS, which they call TopDown18 and that they run themselves. In brief, they run TopDown18 with various parameter settings and, separately, sample redistricting plans using various map drawing methods. Then they analyze the scale and characteristics of the noise from TopDown18, aggregated up to the sampled districts, for the parameters and map drawing methods tested.
One limitation of this approach is that TopDown18’s algorithm, implementation, and parameters are very different than the final version of the 2020 DAS. The other important limitation is the need to resort to using less-than-perfect data as input to the DAS, the real data being confidential. Cohen et al. used a reconstructed version of the 2010 Census data constructed to be consistent with the official tabulations published in SWAP [12]. There is no way to know how the reconstructed dataset differs from the real data, or how those differences might have affected the analyses in [12].
@@ 5 Balancing District Populations with Noise
This section investigates how perturbations from disclosure avoidance interact with population balance requirements. We consider a data user’s ability to use the public DEMO dataset (after disclosure avoidance) to generate redistricting plans satisfying a maximum population deviation target under the confidential CEF dataset (before disclosure avoidance). Because SWAP and CEF agree on population totals, we can use SWAP as a perfect stand-in for CEF.
A significant fraction of redistricting plans satisfying a given population deviation limit on DEMO exceeds that limit on SWAP, as first shown by Kenny et al. [24]. They conclude that it is impossible to produce redistricting maps that adhere to One Person, One Vote requirements on the unseen SWAP dataset, pointing to state legislative districts as particularly challenging.
This section challenges that conclusion. We propose a straightforward method for ensuring that state legislative maps meet population balance thresholds under both datasets. Our method—requiring a smaller population deviation on the public DEMO dataset—is very effective, and does not greatly increase the difficulty of finding valid plans [^9] As evidence of the fact that decreasing this threshold does not greatly increase the difficulty of finding valid plans, all of the ensembles analyzed in Section 6 are sampled using population tolerance offsets derived in Section 5. nor require the adoption of new redistricting techniques. See Section 7 for a discussion of limitations.
We focus on state legislative districts for a few reasons. First, some population deviation is allowed but precise thresholds are still observed. Our techniques would be ill-suited for studying Congressional maps, where population deviations greater than 1 person must be justified, because our algorithms are unable to generate ensembles of maps such strict requirements. Second, states offer enough diversity to observe meaningful variation, while also being few enough to cover exhaustively. Third, we are building on the existing case study of the LA state senate in [24].
@@ 5.1 Meeting Population Deviation Limits Using Offsets
We begin with a case study of the Louisiana state senate—where the sort of population discrepancies we study were first observed by Kenny et al. [24, Figure S4.4]—then extend the analysis to 93 state legislative geographies.
@@
We propose a simple approach: draw the plan to satisfy a slightly tighter limit $\tau-\Delta$ on DEMO, for some value $0\leq\Delta\leq\tau$. We call $\Delta$ the offset. To evaluate this approach, we measure the discrepancy at $\tau$ with offset $\Delta$: the fraction of plans that exceed $\tau$ deviation under SWAP among an ensemble of plans with deviation at most $\tau-\Delta$ under DEMO. When $\Delta=0$, we call this the discrepancy at $\tau$ or the no-offset discrepancy rate. Finally, we also consider what we call the critical offset: the smallest offset with 0 discrepancies observed in our ensembles of 100,000 plans (see Appendix A for more information on the computation of the critical offset).
Conceptually, this is an existence exercise: for our district-construction procedure, does there exist an offset $\Delta$ such that plans drawn to satisfy $\tau-\Delta$ on DEMO are empirically likely (in our ensembles) to satisfy $\tau$ on SWAP. The resulting critical offsets are therefore specific to our simulation method and choice of geographic building blocks (e.g., precincts versus blocks), and should not be treated as universal constants for all redistricting processes (Section 7). Still, this approach can provide a useful framework for reasoning about the effects of disclosure avoidance on OPOV concerns.
5.1.1 A Case Study: The Louisiana State Senate
@@
Figure 1: Discrepancy with offsets in the Louisiana state senate. We plot the fraction of plans exceeding intended population tolerance limit $\tau$, with various offsets $\Delta$ for the Louisiana state senate (i.e., discrepancy at $\tau$ with offset $\Delta$). Dots (solid lines) are computed using the DEMO and SWAP datasets with ensembles of 100,000 plans for each $\tau$ and $\Delta$ (see Sec. 5.1). Squares (dashed lines) are computed from 100,000 samples from the statistical model of district populations and disclosure avoidance noise described in Sec. 5.2. Black triangles (dotted lines) represent the results obtained by [24, Figure S4.4].
Replicating [24], we measure the no-offset discrepancy for each of six population deviation thresholds $\tau$ between 0.1% and 30%. Namely, we generate an ensemble of 100,000 plans for the LA state senate with population deviation at most $\tau$ according to DEMO, and count what fraction of those plans exceed $\tau$ deviation according to SWAP.
@@
The results are summarized by the solid blue line in Figure 1. The overall conclusion remains the same as in [24]: many plans drawn to a satisfy $\tau$ population deviation using DEMO exceed that threshold on SWAP. The exact numbers differ, perhaps due to differences in our experimental setups, such as a different MCMC sampler implementation and larger ensembles.
To test the effectiveness of offsets as a mitigation, we measure the discrepancy at $\tau$ and varying the offset parameter $\Delta$, for the same six values of $\tau$ as before. Additionally, we compute the critical offset for the Louisiana State Senate for $\tau=5\%$ to be $\Delta=0.33\%$ (averaged over 10 runs).
@@
The solid lines in Figure 1 illustrate the results for $\Delta=0.05\%$, 0.1%, 0.2%, and 0.5%. Using $\Delta=0.1\%$ reduces the discrepancy at $\tau=5\%$ from 23.4% to 4.0%. Using $\Delta=0.2\%$ reduces the discrepancy at $\tau=1\%$ from 74.2% to 1.1%. Notably, the discrepancy at $\tau=5\%$ with offset $\Delta=0.5\%$ was 0%. That is, of 100,000 plans generated using a 4.5% tolerance on DEMO, none exceeded a 5% tolerance on SWAP.